OpenStudy (anonymous):

How would you turn this into an equivalent factored form: 14x+18x^2

OpenStudy (cwrw238):

2x(9x + 7)

OpenStudy (anonymous):

How did you do that? Steps.

OpenStudy (accessdenied):

you find common factors between all of the terms to take out by 'reverse'-distributive property first both terms have at least one x and are divisible by 2, so those would be the common factors between them you can re-distribute the numbers back in again to see that it is equal

OpenStudy (cwrw238):

greatest common factor = 2x divide 2x into 14x + 18x^2 gives you whats in the parentheses

OpenStudy (anonymous):

14x+8x^2 Find common coefficient factors 2(7x+4x^2)=(2*(7x)+2(2x^2)) Find common variable factors 2x(7+4x) The laws of exponents show that when we divide common factors division is treated as subtraction \[x^2/x =x^{2-1}=x\]

OpenStudy (anonymous):

Thanks guys.

OpenStudy (anonymous):

18x^2 + 14x = 2x ( 9x + 7)